OpenStudy (anonymous):

Explain, in complete sentences, why the expression x2 + 28x + 60 is prime.

OpenStudy (anonymous):

@Hero @amistre64

OpenStudy (amistre64):

what do we call things that are prime? think of a prime number, what property does it hold that makes it prime?

OpenStudy (yoyo123):

Because you can't factor it, in order to factor that polynomial you'd have to find 2 numbers that multiply to 60 and add to 28, which there are none.

OpenStudy (amistre64):

if you know about determinants, then you will be able to assess if its prime or not

OpenStudy (tkhunny):

There ain't no factors of 60 that sum to 28. Is that a sentence?

OpenStudy (amistre64):

might be discriminants .... to many 'd' words to keep track of :)

OpenStudy (anonymous):

@amistre64 a prime number is 5

OpenStudy (amistre64):

irrational roots are still factors tho

OpenStudy (amistre64):

why is 5 prime?

OpenStudy (anonymous):

because it can only be divided evenly by 1 or 5

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

right, it cant be factored into smaller parts. with polymonials, they are prime if they have no 'real' valued roots .... if they possess complex roots, they are not factorable across the reals

OpenStudy (amistre64):

your poly is prime if and only if it has no 'real' roots. and that can be determined by the discriminant: b^2 -4ac = D if D >=0 it has real roots if D < 0 it has no real roots

OpenStudy (amistre64):

take x^2+4 as an example: (x+2i)(x-2i) = x^2 -(2i)^2 = x^2 + 4, it has factors that are complex, not real. 0^2 -4(1)(1) = -4, -4<0 therefore x^2 + 4 is prime

OpenStudy (anonymous):

Thank You for helping me @amistre64

OpenStudy (amistre64):

youre welcome